OpenStudy (anonymous):

The following is an indirect proof of the Addition Property of Equality: For real numbers a, b, and c, if a = b, then a + c = b + c. Assume____________. According to the given information, a = b. By the Subtraction Property of Equality, one can subtract the same number from both sides of an equation without changing the equation. Therefore, a + c - c ≠ b + c - c. Through subtraction, the c's cancel and a ≠ b. This contradicts the given information so a + c = b + c.

OpenStudy (anonymous):

Which statement accurately completes the proof? Select one: a. a + c = b + c b. a + c ≠ b + c c. a ≠ b d. b = c

OpenStudy (anonymous):

@Hero @thomaster

OpenStudy (anonymous):

by reading the question I think the answer is B idk how to explain why tho sorry.... if I helped give me a best response plz

OpenStudy (anonymous):

hey wait dont go

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

here?

OpenStudy (anonymous):

lets talk in inbox

OpenStudy (anonymous):

first look at my question and see if u can answer my question!! LOL.. http://openstudy.com/study#/updates/5250385de4b0bfb1fb43e775

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ewww you changed ur profile pic... the other one was cuter btw profile pic thing

OpenStudy (anonymous):

im going to my question now